OpenStudy (anonymous):

Enough of a monoprotic acid is dissolved in water to produce a 0.0136 M solution. The pH of the resulting solution is 2.45. Calculate the Ka for the acid.

OpenStudy (abb0t):

Find the concentration first: [H+]: \[10^{-2.45}\]

OpenStudy (anonymous):

then???

OpenStudy (abb0t):

You know that for any monoprotic acid, it dissociates in the following manner: \[HA \rightleftharpoons H^+ + A^-\]

OpenStudy (abb0t):

Are you familiar with how to make an ICE table?

OpenStudy (anonymous):

[H+] = Ka - \[[H^+] = Ka - \frac{ \log [A] }{\log [B] ? }\]

OpenStudy (anonymous):

lol is the equation right? =))

OpenStudy (anonymous):

then find the -log of everything

OpenStudy (anonymous):

it results into this equation: \[pH = pKa + \frac{ \log [B] }{ \log [A]? }\] wherein A is the weak acid and B is the conjugate base

OpenStudy (anonymous):

to find the Ka, you have to derive the formula

OpenStudy (anonymous):

im not sure how to set up the ICE table..

OpenStudy (anonymous):

uh... i think that the info needed is incomplete? i'm not sure though O.O

OpenStudy (anonymous):

where did you get that question by the way? :))

OpenStudy (anonymous):

saplinglearning.com

OpenStudy (anonymous):

oh, okay. :)) ugh, i need to review Henderson-Hassellbalch equation @_@